Can someone please help me solve the following problem:
What would be the y-intercept of the equation: y=-3/2x^2+5x-6
THANK YOU!!


Answer by nerdybill(7384)

To find the y-intercept, simply set 'x' to 0 and solve:
y=-3/2x^2+5x-6
y=-3/2(0)^2+5(0)-6
y=0+0-6
y = -6
.
therefore, the y-intercept is at (0,-6)